### v1

# Notes on Bi-parameter Paraproducts

Abstract.
In this note, we investigate the sharpness of existing bounds for various types of bi-parameter paraproducts acting between product Hardy spaces in the dyadic setting. We show that these bounds are sharp in most cases but fail to be so in one particular instance.
Key words and phrases:
Hardy spaces, Paraproducts
1991 Mathematics Subject Classification:
42B30, 42B35
1. Introduction
This work is a sequel to our previous two papers on the boundedness properties of paraproducts [20, 21]. Here, we are concerned with sharpness of the existing bounds for the remaining dyadic bi-parameter paraproducts not studied in [20]. For various applications and properties of paraproducts, we refer the reader to [3, 2, 15, 16, 17, 14].
Let us begin by recalling the one-parameter dyadic paraproducts, three bilinear forms defined by
$\pi^{\varepsilon}(f,g)(x):=\sum_{I\in\mathcal{D}}\left\langle f,h_{I}^{\varepsilon_{1}}\right\rangle\left\langle g,h_{I}^{\varepsilon_{2}}\right\rangle h_{I}^{\varepsilon_{1}+\varepsilon_{2}}(x),\quad\varepsilon=(\varepsilon_{1},\varepsilon_{2})\in\{0,1\}^{2}\backslash\{(0,0)\},\quad x\in\mathbb{R},$
where $I$ stands for a dyadic interval on the line, $\mathcal{D}$ denotes the collection of all such intervals, $h_{I}$ is the Haar wavelet associated with $I$, and $h_{I}^{\epsilon}$ is given by
$h_{I}^{\epsilon}(x):=\begin{cases}h_{I}(x),&\epsilon=1,\\ \bar{\chi}_{I}(x),&\epsilon=0,\end{cases}\quad x\in\mathbb{R},$
where $\bar{\chi}_{I}:=\frac{\chi_{I}}{|I|}$, and the sum $\varepsilon_{1}+\varepsilon_{2}$ is understood modulo 2. It is then easy to see that, for sufficiently nice functions, we have
[…]
To avoid unnecessary complications, throughout this note we make the qualitative assumption that all our functions, on either $\mathbb{R}$ or $\mathbb{R}^{2}$, are real-valued and simple, in the sense that they are finite linear combinations of characteristic functions of dyadic intervals or squares. Since none of the bounds depend on this a priori assumption, standard limiting arguments extend our results to the general case. See also [20, 21], where a different approach has been taken.
Fairly well-known arguments show that, for $0<p<\infty$, the bilinear form $\pi^{(0,1)}$ is bounded from $H^{p}_{d}(\mathbb{R})\times BMO_{d}(\mathbb{R})$ and $L^{\infty}(\mathbb{R})\times H^{p}_{d}(\mathbb{R})$ to $H^{p}_{d}(\mathbb{R})$. Here, for $0<p<\infty$, $H^{p}_{d}(\mathbb{R})$ denotes the dyadic Hardy space on the line, (quasi-)normed by
$\|f\|_{H^{p}_{d}(\mathbb{R})}:=\|M(f)\|_{L^{p}(\mathbb{R})},\quad M(f)(x):=\sup_{x\in I}|\left\langle f\right\rangle_{I}|,\quad x\in\mathbb{R},$
where we use $\left\langle f\right\rangle_{E}$ for the average of $f$ over a measurable set $E$. Also, $BMO_{d}(\mathbb{R})$ denotes the space of functions with uniformly bounded mean oscillation on dyadic intervals, i.e.
$\|f\|_{BMO_{d}(\mathbb{R})}:=\sup_{I\in\mathcal{D}}\big(\frac{1}{|I|}\sum_{I^{\prime}\subseteq I}\left\langle f,h_{I^{\prime}}\right\rangle^{2}\big)^{\frac{1}{2}}<\infty.$
We refer the reader to [22] for more on these spaces.
Because of the symmetry $\pi^{(0,1)}(f,g)=\pi^{(1,0)}(g,f)$, a similar result holds for $\pi^{(1,0)}$. In addition, both forms are bounded from $H^{p}_{d}(\mathbb{R})\times H^{r}_{d}(\mathbb{R})$ to $H^{q}_{d}(\mathbb{R})$, with $\frac{1}{q}=\frac{1}{p}+\frac{1}{r}$ and $0<r,p,q<\infty$. These two results also hold for $\pi^{(1,1)}$, but only in the reflexive range of exponents, i.e., where $1<r,p,q<\infty$. Regarding the sharpness of these boundedness properties, a natural approach is to freeze one of the inputs and study the operator norm of the resulting linear operators. Because of the symmetries, it is enough to study the linear operators
(2) $\pi_{g}(f)=\sum_{I\in\mathcal{D}}\left\langle f\right\rangle_{I}\left\langle g,h_{I}\right\rangle h_{I},\quad\pi^{\prime}_{g}(f)=\sum_{I\in\mathcal{D}}\left\langle f,h_{I}\right\rangle\left\langle g,h_{I}\right\rangle\bar{\chi}_{I},\quad\pi^{\prime\prime}_{g}(f)=\sum_{I\in\mathcal{D}}\left\langle f,h_{I}\right\rangle\left\langle g\right\rangle_{I}h_{I}.$
We may also exclude $\pi^{\prime}_{g}$ from our study, since it is the adjoint of $\pi_{g}$ and therefore has the same norm in the Banach range of exponents. In [4], it was shown that
$\|\pi_{g}\|_{L^{p}(\mathbb{R})\to L^{p}(\mathbb{R})}\simeq\|g\|_{BMO_{d}(\mathbb{R})},\quad 1<p<\infty,$
and in [11], the authors established the equivalence
$\|\pi_{g}\|_{L^{p}(\mathbb{R})\to L^{q}(\mathbb{R})}\simeq\|g\|_{L^{r}(\mathbb{R})},\quad\frac{1}{q}=\frac{1}{p}+\frac{1}{r},\quad 1<p,r,q<\infty.$
Recently, in [21], we showed that both results extend to the full range of exponents, provided that Lebesgue spaces are replaced with dyadic Hardy spaces. The operator $\pi^{\prime\prime}_{g}$ is simpler to deal with, and one can easily verify that
$\|\pi^{\prime\prime}_{g}\|_{H_{d}^{p}(\mathbb{R})\to H_{d}^{p}(\mathbb{R})}\simeq\|g\|_{L^{\infty}(\mathbb{R})},\quad 0<p<\infty.$
It is enough to recall the square function characterization of dyadic Hardy spaces, which states that
$\|f\|_{H_{d}^{p}(\mathbb{R})}\simeq\|S(f)\|_{L^{p}(\mathbb{R})},\quad 0<p<\infty,\quad S(f):=\big(\sum_{I\in\mathcal{D}}\left\langle f,h_{I}\right\rangle^{2}\bar{\chi}_{I}\big)^{\frac{1}{2}}.$
As expected, the $H_{d}^{p}(\mathbb{R})$-to-$H_{d}^{q}(\mathbb{R})$ norm of $\pi^{\prime\prime}_{g}$ is also comparable to $\|g\|_{H_{d}^{r}(\mathbb{R})}$ in the full range of exponents, when $\frac{1}{q}=\frac{1}{p}+\frac{1}{r}$. In the reflexive range, this result was established in [18] (Theorem 12.2, p. 128), and in Theorem 2.1 we give a new proof in the product setting, to which we now turn.
Let us recall the family of bi-parameter paraproducts: the nine bilinear forms arising from the product of two functions on the plane, expanded in the rectangular Haar basis. Let
$h_{I}\otimes h_{J}(x,y)=h_{I}(x)h_{J}(y),\quad(x,y)\in\mathbb{R}^{2},$
[…]
(3) $\pi^{\varepsilon}(f,g):=\pi^{\varepsilon_{1}}\otimes\pi^{\varepsilon_{2}}(f,g)=\sum_{I,J\in\mathcal{D}}\left\langle f,h_{I}^{\varepsilon_{11}}\otimes h_{J}^{\varepsilon_{21}}\right\rangle\left\langle g,h_{I}^{\varepsilon_{12}}\otimes h_{J}^{\varepsilon_{22}}\right\rangle h_{I}^{\varepsilon_{11}+\varepsilon_{12}}\otimes h_{J}^{\varepsilon_{21}+\varepsilon_{22}}.$
Similarly to (1), we have
$f(x,y)g(x,y)=\sum_{\begin{subarray}{c}\varepsilon=(\varepsilon_{1},\varepsilon_{2})\\ \varepsilon_{1},\varepsilon_{2}\neq(0,0)\end{subarray}}\pi^{\varepsilon}(f,g)(x,y),\quad(x,y)\in\mathbb{R}^{2},$
which directly follows from (1) when $f=f_{1}\otimes f_{2}$ and $g=g_{1}\otimes g_{2}$, and from linearity in the general case. When $\varepsilon_{1}=\varepsilon_{2}$, the bilinear form $\pi^{\varepsilon}$ is called an “unmixed” paraproduct, since in such cases the cancellative terms $h_{I},h_{J}$ and the non-cancellative terms $\bar{\chi}_{I},\bar{\chi}_{J}$ are separated from each other in all three terms of the sum in (3). The other forms are referred to as “mixed” paraproducts.
To discuss boundedness of these forms and the sharpness of the existing results, we take a similar approach as explained before. Considering the symmetries in $f$ and $g$, and in $x$ and $y$, and excluding the adjoint operators, one identifies four different linear operators, listed below.
$\pi^{1}_{g}(f):=\pi^{(0,1)}\otimes\pi^{(0,1)}(f,g)=\sum_{I,J\in\mathcal{D}}\left\langle f,\bar{\chi}_{I}\otimes\bar{\chi}_{J}\right\rangle\left\langle g,h_{I}\otimes h_{J}\right\rangle h_{I}\otimes h_{J},$
$\pi^{2}_{g}(f):=\pi^{(1,0)}\otimes\pi^{(1,0)}(f,g)=\sum_{I,J\in\mathcal{D}}\left\langle f,h_{I}\otimes h_{J}\right\rangle\left\langle g,\bar{\chi}_{I}\otimes\bar{\chi}_{J}\right\rangle h_{I}\otimes h_{J}$
$\pi^{3}_{g}(f):=\pi^{(0,1)}\otimes\pi^{(1,0)}(f,g)=\sum_{I,J\in\mathcal{D}}\left\langle f,\bar{\chi}_{I}\otimes h_{J}\right\rangle\left\langle g,h_{I}\otimes\bar{\chi}_{J}\right\rangle h_{I}\otimes h_{J}$
$\pi^{4}_{g}(f)=\pi^{(0,1)}\otimes\pi^{(1,1)}(f,g)=\sum_{I,J\in\mathcal{D}}\left\langle f,\bar{\chi}_{I}\otimes h_{J}\right\rangle\left\langle g,h_{I}\otimes h_{J}\right\rangle h_{I}\otimes\bar{\chi}_{J}.$
In the rest of this note, we discuss the boundedness of the above linear operators acting between different product Hardy spaces. We refer the reader to Theorem 2.1, Theorem 3.2, and the example presented at the end of Section 3 for a quick overview of the new results of this note.
Before proceeding, it is convenient to simplify the notation. In $\mathbb{R}^{2}$, we use $I$ to denote a dyadic interval on the $x$-axis and $J$ for such intervals on the $y$-axis. From now on, we write $R=I\times J$ for a dyadic rectangle in the plane, set $h_{R}=h_{I}\otimes h_{J}$, and for a function $f$, define $f_{R}=\langle f,h_{R}\rangle$. In addition, for a function $f$ on $\mathbb{R}^{2}$, we let
$f_{I}(y):=\langle f(\cdot,y),h_{I}\rangle,\qquad f_{J}(x):=\langle f(x,\cdot),h_{J}\rangle.$
2. Unmixed Paraproducts
Let us begin with the operator $\pi^{1}_{g}$, which was the main object of our recent work [20], and is given by
$\pi^{1}_{g}(f):=\sum_{R}\left\langle f\right\rangle_{R}g_{R}h_{R}.$
There, we have shown that the behavior of this operator and its one-parameter analogue $\pi_{g}$, defined in (2), are identical. More precisely, we have shown that
$\|\pi^{1}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\simeq\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})},\quad 0<p<\infty,$
$\|\pi^{1}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{q}(\mathbb{R}\otimes\mathbb{R})}\simeq\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})},\quad\frac{1}{q}=\frac{1}{p}+\frac{1}{r},\quad 0<p,r,q<\infty,$
where in the above $H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})$ stands for the bi-parameter dyadic Hardy space, (quasi-)normed by
$\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}:=\|M(f)\|_{L^{p}(\mathbb{R}^{2})},\quad M(f)(x,y):=\sup_{(x,y)\in R}|\left\langle f\right\rangle_{R}|,\quad(x,y)\in\mathbb{R}^{2},$
and $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$ denotes the dyadic product BMO, i.e., the space of functions with
$\|f\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}:=\sup_{\Omega}\big(\frac{1}{|\Omega|}\sum_{R\subseteq\Omega}f_{R}^{2}\big)^{\frac{1}{2}}<\infty,$
where the supremum is taken over all open sets of finite positive measure. Since it will be clear from the context, we use the same notation $M$ for the dyadic maximal operator on the line and the strong operator defined above. We also refer the reader to [20] for the reason behind the above equivalences, and to [22, 6, 9, 7] for more on Hardy spaces in the product setting. See also [5, 19]. The adjoint of this operator, which is unmixed too, is given by
$(\pi^{1}_{g})^{t}(f)=\pi^{(1,1)}\otimes\pi^{(1,1)}(f,g)=\sum_{R}f_{R}g_{R}\frac{\chi_{R}}{|R|},$
and therefore it satisfies similar properties, but only in the Banach range of spaces.
Since $\pi^{1}_{g}$ is well-understood, we turn to the second operator, $\pi^{2}_{g}$, which is simply defined by
$\pi^{2}_{g}(f):=\sum_{R}f_{R}\left\langle g\right\rangle_{R}h_{R},$
and satisfies the known bounds
(4) $\|\pi^{2}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\lesssim\|g\|_{L^{\infty}(\mathbb{R}^{2})},\quad 0<p<\infty,$
(5) $\|\pi^{2}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{q}(\mathbb{R}\otimes\mathbb{R})}\lesssim\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})},\quad\frac{1}{q}=\frac{1}{p}+\frac{1}{r},\quad 0<p,r,q<\infty,$
which can be easily verified after recalling the square function characterization of $H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})$, i.e., the fact that
$\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\simeq\|S(f)\|_{L^{p}(\mathbb{R}^{2})},\quad 0<p<\infty,\quad S(f):=\Big(\sum_{R}|f_{R}|^{2}\frac{\chi_{R}}{|R|}\Big)^{\frac{1}{2}}.$
Now, the above bounds are simple consequences of the crucial pointwise inequality
$S(\pi^{2}_{g}(f))\leq S(f)M(g),$
the square function characterization of product Hardy spaces, and Hölder’s inequality.
Regarding the sharpness of (4) and (5), we could not find anything in the literature. Thus, as our first theorem, we show that, just like in the one-parameter setting, both of these inequalities are indeed equivalences. Before doing so, let us briefly recall the notion of Carleson families, the John-Nirenberg lemma in the product setting, and a weak form of the Fefferman-Cordoba covering lemma for rectangles.
A collection of rectangles (with sides parallel to the axes), $\mathcal{C}$, is called $\Lambda$-Carleson if, for every open set $\Omega$, there holds
$\sum_{\begin{subarray}{c}R\subseteq\Omega\\ R\in\mathcal{C}\end{subarray}}|R|\leq\Lambda|\Omega|.$
As is well-known, this condition is equivalent to $\eta$-sparseness with $\eta=\Lambda^{-1}$ [10, 12]. The sparseness condition means that every element $R$ of the family has a piece $E_{R}\subset R$ of density at least $\eta$ such that all these pieces are pairwise disjoint. Then the John-Nirenberg lemma states that an $\eta$-sparse family of rectangles is essentially a disjoint family in the sense that
$\|\sum_{\begin{subarray}{c}R\subseteq\Omega\\ R\in\mathcal{C}\end{subarray}}\chi_{R}\|_{L^{p}(\mathbb{R}^{2})}\lesssim_{p,\eta}|\Omega|^{\frac{1}{p}},\quad 0<p<\infty.$
Finally, it follows from the Fefferman-Cordoba covering lemma that, for any collection of rectangles $\mathcal{C}$, it is possible to extract a $\frac{1}{2}$-sparse sub-collection $\mathcal{C}^{\prime}$ such that
$|\cup_{R\in\mathcal{C}}R|\simeq|\cup_{R^{\prime}\in\mathcal{C}^{\prime}}R^{\prime}|.$
See [9, 17] for a detailed exposition of these.
Theorem 2.1.
For any function $g$, both bounds in (4) and (5) are indeed equivalences.
Proof.
The first case follows by plugging $f=h_{R}$ into the operator and using Lebesgue’s differentiation theorem. We now show that (5) is sharp. To this aim, normalize $g$ such that
$\|\pi^{2}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{q}(\mathbb{R}\otimes\mathbb{R})}=1.$
Then we need to show that $\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})}\lesssim 1$, and to do so, we build a proper test function by looking at the level sets of $M(g)$.
For an integer $k$, observe that the level set $\{M(g)>2^{k}\}$ is a union of rectangles $R$ with the property that $|\left\langle g\right\rangle_{R}|>2^{k}$. Apply the above-mentioned covering lemma to these rectangles and extract a $\frac{1}{2}$-sparse sub-collection $\mathcal{C}_{k}$, whose total measure is proportional to $|\{M(g)>2^{k}\}|$. Then consider $\mathcal{C}=\cup_{k\in\mathbb{Z}}\mathcal{C}_{k}$, and for each rectangle $R\in\mathcal{C}$, let $\lambda(R)$ be the largest $k$ for which $R\in\mathcal{C}_{k}$. Based on our a priori assumption on $g$, this function has compact support and is bounded. Therefore, each $\mathcal{C}$ is finite and $\lambda(R)$ is well-defined. Next, consider the test function
$f=\sum_{R\in\mathcal{C}}2^{t\lambda(R)}|R|^{\frac{1}{2}}h_{R},\quad t=\frac{r}{p},$
satisfying
(6) $\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\lesssim\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})}^{\frac{r}{p}}.$
To see this, we estimate the $L^{p}(\mathbb{R}^{2})$-norm of $S(f)$, and in doing so we consider two separate cases. First, suppose $0<p\leq 2$, in which case we take advantage of sub-linearity and obtain
$S(f)^{p}\leq\sum_{k\in\mathbb{Z}}2^{ptk}\big(\sum_{R\in\mathcal{C}_{k}}\chi_{R}\big)^{\frac{p}{2}},$
which, after integrating and applying the John-Nirenberg lemma, gives
$\|S(f)\|_{L^{p}(\mathbb{R}^{2})}^{p}\leq\sum_{k\in\mathbb{Z}}2^{ptk}\int\big(\sum_{R\in\mathcal{C}_{k}}\chi_{R}\big)^{\frac{p}{2}}\lesssim\sum_{k\in\mathbb{Z}}2^{ptk}\big|\cup_{R\in\mathcal{C}_{k}}R\big|.$
Now, recall that $|\cup_{R\in\mathcal{C}_{k}}R|\simeq|\{M(g)>2^{k}\}|$, and apply the layer-cake formula to obtain
$\|S(f)\|_{L^{p}(\mathbb{R}^{2})}^{p}\lesssim\sum_{k\in\mathbb{Z}}2^{rk}|\{M(g)>2^{k}\}|\simeq\|M(g)\|_{L^{r}(\mathbb{R}^{2})}^{r},$
which is the claimed inequality. Next, consider the case $2<p<\infty$, in which we appeal to duality. Let $\varphi$ be a function with $\|\varphi\|_{L^{(\frac{p}{2})^{\prime}}(\mathbb{R}^{2})}=1$, such that
$\|S(f)\|_{L^{p}}^{2}=\|S(f)^{2}\|_{L^{\frac{p}{2}}}=\int S(f)^{2}\varphi,$
where, as usual, $(\frac{p}{2})^{\prime}$ is the Hölder conjugate of $\frac{p}{2}$. Then we have
$\|S(f)\|_{L^{p}}^{2}=\int\sum_{R\in\mathcal{C}}2^{2t\lambda(R)}\chi_{R}\varphi=\int\sum_{R\in\mathcal{C}}2^{2t\lambda(R)}\left\langle\varphi\right\rangle_{R}|R|\leq\sum_{k\in\mathbb{Z}}2^{2tk}\sum_{R\in\mathcal{C}_{k}}|\left\langle\varphi\right\rangle_{R}||R|.$
At this point, recall that each collection $\mathcal{C}_{k}$ is $\frac{1}{2}$-sparse, and thus for each rectangle $R\in\mathcal{C}_{k}$, we may find a subset $E_{R}\subset R$ with $|R|\leq 2|E_{R}|$, such that the sets $E_{R}$ are disjoint. Therefore, we may continue to estimate the last term by
$\|S(f)\|_{L^{p}}^{2}\lesssim\sum_{k\in\mathbb{Z}}2^{2tk}\sum_{R\in\mathcal{C}_{k}}|\left\langle\varphi\right\rangle_{R}||E_{R}|\leq\sum_{k\in\mathbb{Z}}2^{2tk}\int_{\cup_{R\in\mathcal{C}_{k}}E_{R}}M(\varphi)\leq\sum_{k\in\mathbb{Z}}2^{2tk}\int_{\{M(g)>2^{k}\}}M(\varphi),$
which implies that
$\|S(f)\|_{L^{p}}^{2}\lesssim\int M(\varphi)M(g)^{2t}\leq\|M(\varphi)\|_{L^{(\frac{p}{2})^{\prime}}}\|M(g)\|_{L^{r}}^{\frac{2r}{p}}\lesssim\|M(g)\|_{L^{r}}^{\frac{2r}{p}},$
again yielding the claimed inequality. In the above, we used the boundedness of the strong maximal operator $M$ on $L^{\frac{p}{2}}$, which is allowed since $\frac{p}{2}>1$.
Next, we apply the operator $\pi^{2}_{g}$ to $f$ and obtain
$F=\pi^{2}_{g}(f)=\sum_{R\in\mathcal{C}}2^{t\lambda(R)}\left\langle g\right\rangle_{R}|R|^{\frac{1}{2}}h_{R},$
which has two properties. First,
$S(F)\geq 2^{t\lambda(R)}|\left\langle g\right\rangle_{R}|\chi_{R},\quad R\in\mathcal{C},$
implying that
(7) $S(F)(x,y)>2^{(1+t)k},\quad(x,y)\in\cup_{R\in\mathcal{C}_{k}}R,\quad k\in\mathbb{Z}.$
Second, since $F=\pi^{2}_{g}(f)$ and we have established that $\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\lesssim\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})}^{\frac{r}{p}}$, we must have
(8) $\|S(F)\|_{L^{q}}\lesssim\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})}^{\frac{r}{p}}.$
Therefore, from (7) and the layer-cake formula, we obtain
$\int M(g)^{r}\simeq\sum_{k\in\mathbb{Z}}2^{rk}|\{M(g)>2^{k}\}|\simeq\sum_{k\in\mathbb{Z}}2^{rk}|\cup_{R\in\mathcal{C}_{k}}R|\leq\sum_{k\in\mathbb{Z}}2^{rk}|\{S(F)>2^{(1+t)k}\}|,$
which, after recalling that $1+t=\frac{r}{q}$, implies
$\int M(g)^{r}\lesssim\int S(F)^{q}.$
Finally, combining this with (8) yields
$\|M(g)\|_{L^{r}(\mathbb{R}^{2})}^{\frac{r}{q}}\lesssim\|S(F)\|_{L^{r}(\mathbb{R}^{2})}^{\frac{r}{p}},$
which is exactly $\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})}\lesssim 1$, the desired result. The proof is now complete. ∎
3. Mixed Paraproducts
Now we turn our attention to the last two operators, $\pi^{3}_{g}$ and $\pi^{4}_{g}$, which are of mixed type. The analysis of such operators naturally leads one to consider the mixed square-maximal or maximal-square operators of the form
$S_{2}M_{1}(f)(x,y):=\Big(\sum_{J\in\mathcal{D}}M(f_{J})^{2}(x)\bar{\chi}_{J}(y)\Big)^{\frac{1}{2}},\quad M_{1}S_{2}(f)(x,y):=\sup_{x\in I}\Big(\sum_{J\in\mathcal{D}}|\left\langle f_{J}\right\rangle_{I}|^{2}\bar{\chi}_{J}(y)\Big)^{\frac{1}{2}},$
and, quite similarly,
$S_{1}M_{2}(g)(x,y):=\Big(\sum_{I\in\mathcal{D}}M(g_{I})^{2}(y)\bar{\chi}_{I}(x)\Big)^{\frac{1}{2}},\quad M_{2}S_{1}(g)(x,y):=\sup_{y\in J}\Big(\sum_{I\in\mathcal{D}}|\left\langle g_{I}\right\rangle_{J}|^{2}\bar{\chi}_{I}(x)\Big)^{\frac{1}{2}}.$
It is then not hard to see, and it is well-known, that each of these mixed operators gives another characterization of $H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})$ [15, 16, 14, 17]. This means that for $0<p<\infty$, the $L^{p}(\mathbb{R}^{2})$-norm of each of the four functions above is comparable to the $H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})$-norm of $f$ and $g$, respectively. Let us briefly sketch the proof of these equivalences, say for $f$. First, note that the pointwise inequality
$M_{1}S_{2}(f)\leq S_{2}M_{1}(f),$
implies
$\|M_{1}S_{2}(f)\|_{L^{p}(\mathbb{R}^{2})}\leq\|S_{2}M_{1}(f)\|_{L^{p}(\mathbb{R}^{2})}.$
To see why
$\|S_{2}M_{1}(f)\|_{L^{p}(\mathbb{R}^{2})}\lesssim\|S(f)\|_{L^{p}(\mathbb{R}^{2})},$
fix $y$ and observe that the vector-valued Fefferman-Stein inequality for the maximal operator on the line, $M$, together with the square function characterization of $L^{p}(\mathbb{R})$, implies that the operator $S_{2}M_{1}$ is bounded on all $L^{p}(\mathbb{R})$ with $1<p<\infty$. This, combined with Fubini’s theorem, establishes the above inequality for $1<p<\infty$. The case $0<p\leq 1$ then follows from the atomic decomposition of $H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})$. To complete the chain of equivalences, one has to show that
$\|S(f)\|_{L^{p}(\mathbb{R}^{2})}\lesssim\|M_{1}S_{2}(f)\|_{L^{p}(\mathbb{R}^{2})},$
which, after using Fubini’s theorem and fixing $y$ again, follows from the equivalence of the square function and maximal characterizations of $H_{d}^{p}(\mathbb{R},l^{2})$.
3.1. The Operator $\pi^{3}_{g}$
Now, we continue with the boundedness properties of the first mixed paraproduct, $\pi^{3}_{g}$, given by
$\pi^{3}_{g}(f):=\sum_{I,J\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}\left\langle g_{I}\right\rangle_{J}h_{I}\otimes h_{J}.$
For this operator, just like for $\pi^{1}_{g}$ and $\pi^{2}_{g}$, there holds that
(9) $\|\pi^{3}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{q}(\mathbb{R}\otimes\mathbb{R})}\lesssim\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})},\quad\frac{1}{q}=\frac{1}{p}+\frac{1}{r},\quad 0<p,r,q<\infty,$
which is a simple consequence of the pointwise inequality
$S(\pi^{3}_{g}(f))\leq S_{2}M_{1}(f)M_{2}S_{1}(g),$
and the above-mentioned mixed characterizations of product Hardy spaces [15, 16, 14, 17].
Unfortunately, we were unable to establish the sharpness of (9) in the full range of exponents. This is because the structure of the level sets of the involved mixed operators and their relationships is more complicated, and the picture is obscured. However, in Theorem 3.2, we show that when $1<r<\infty$, the lower bound follows directly from the one for the operator $\pi_{g}$ defined in (2).
Regarding the diagonal counterpart of (9), we show that the operator norm is comparable to a mixed-type norm of $g$, i.e.,
$\Big\|\|g(x,y)\|_{BMO_{d}(\mathbb{R},dx)}\Big\|_{L^{\infty}(\mathbb{R},dy)},$
which is simply the essential supremum of the $BMO_{d}(\mathbb{R})$-norm of the horizontal slices of $g$ (a similar notation is used for a mixture of $H^{r}$ and $L^{r}$ norms in (12)). See [15, 14], with the above quantity replaced by the $L^{\infty}(\mathbb{R}^{2})$-norm. We will establish this using the atomic decomposition theorem from [20]. To this aim, let us recall some simple but useful notions.
In the plane, a family of measurable sets $\Omega_{i}$, with $i=0,1,2,\ldots$, is called contracting if
$\Omega_{i+1}\subset\Omega_{i},\quad|\Omega_{i+1}|\leq\frac{1}{2}|\Omega_{i}|,\quad i=0,1,2,\ldots.$
It is then easy to see that the maximal operator associated with such a family,
$m(g)(z):=\sup_{\begin{subarray}{c}z\in\Omega_{i}\\ i\geq 0\end{subarray}}\left\langle|g|\right\rangle_{\Omega_{i}},$
is bounded on $L^{p}(\mathbb{R}^{2})$ for $1<p\leq\infty$.
Next, recall that in the product setting a function $f$ is called an $L^{s}$-atom ($\max(1,p)<s<\infty$) supported on $\Omega$ if
$f=\sum_{R\subseteq\Omega}f_{R}h_{R},\quad\|f\|_{L^{s}(\mathbb{R}^{2})}\leq|\Omega|^{\frac{1}{s}}.$
It is then simple to see that $\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\lesssim|\Omega|^{\frac{1}{p}}$. Finally, for any function $f$ and any $0<p<\infty$, one may find a contracting family of open sets $\{\Omega_{i}\}$ and $L^{s}$-atoms $f_{i}$ supported on $\Omega_{i}$ such that
(10) $f=\sum_{i}a_{i}f_{i},\quad\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\simeq\Big(\sum_{i}a_{i}^{p}|\Omega_{i}|\Big)^{\frac{1}{p}}.$
Also, when $f$ is finite linear combination of rectangular Haar function, the this sum is finite. See [20] for the proof. In order to use this atomic decomposition we need to prove a simple lemma which is very useful when working with “local operators”.
Lemma 3.1.
Let $T$ be a linear operator that is local in the sense that it maps $L^{s}$-atoms supported on $\Omega$ into $L^{q}$-atoms supported on the same set $\Omega$ ($1<s,q<\infty$). Then $T$ is bounded on $H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})$ for $0<p<q$. Moreover, the same conclusion holds for $1<p<q$, provided that for any $L^{s}$-atom $f$ supported on $\Omega$, $T(f)$ is supported on $\Omega$ (not necessarily an atom) and satisfies
$\|T(f)\|_{L^{q}(\mathbb{R}^{2})}\leq|\Omega|^{\frac{1}{q}}.$
Proof.
Take $f\in H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})$, and without loss of generality assume that $f$, has a finite Haar support. Then apply the atomic decomposition (10) and obtain $f_{i}$’s and $\Omega_{i}$’s. We have that
$T(f)=\sum_{i}a_{i}T(f_{i}),$
implying that for $0<p\leq 1$, there holds
$\|T(f)\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}^{p}\leq\sum_{i}a_{i}^{p}\|T(f_{i})\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}^{p}\lesssim\sum_{i}a_{i}^{p}|\Omega_{i}|\lesssim\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}^{p},$
which proves the claim. To treat the case $1<p<\infty$, we appeal to duality and pair $T(f)$ with a function $\varphi$ with $\|\varphi\|_{L^{p^{\prime}}(\mathbb{R}^{2})}=1$. Then
$\left\langle T(f),\varphi\right\rangle=\sum_{i}a_{i}\left\langle T(f_{i}),\varphi\right\rangle\leq\sum_{i}|a_{i}|\left\langle|T(f_{i})|^{q}\right\rangle_{\Omega_{i}}^{\frac{1}{q}}\left\langle|\varphi|^{q^{\prime}}\right\rangle_{\Omega_{i}}^{\frac{1}{q^{\prime}}}|\Omega_{i}|\leq\sum_{i}|a_{i}|\left\langle|\varphi|^{q^{\prime}}\right\rangle_{\Omega_{i}}^{\frac{1}{q^{\prime}}}|\Omega_{i}|\leq 2$
$\sum_{i}|a_{i}|\left\langle|\varphi|^{q^{\prime}}\right\rangle_{\Omega_{i}}^{\frac{1}{q^{\prime}}}|\Omega_{i}\backslash\Omega_{i+1}|\leq\int m(|\varphi|^{q^{\prime}})^{\frac{1}{q^{\prime}}}\sum_{i}|a_{i}|\chi_{\Omega_{i}\backslash\Omega_{i+1}}\lesssim\big(\sum_{i}a_{i}^{p}|\Omega_{i}|\big)^{\frac{1}{p}}\lesssim\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})},$
proving the claim. Note that in the last line we used Hölder’s inequality, boundedness of the operator $m$ on $L^{\frac{p^{\prime}}{q^{\prime}}}$, with $q^{\prime}<p^{\prime}$, as well as disjointness of the sets $\Omega_{i}\backslash\Omega_{i+1}$. The proof is now complete. ∎
Theorem 3.2.
For any function $g$, there holds
(11) $\|\pi^{3}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\simeq\left\|\|g(x,y)\|_{BMO_{d}(\mathbb{R},dx)}\right\|_{L^{\infty}(\mathbb{R},dy)},\quad 0<p<\infty.$
Also, for $0<p,r,q<\infty$ with $\frac{1}{q}=\frac{1}{p}+\frac{1}{r}$, we have
(12) $\left\|\|g(x,y)\|_{H_{d}^{r}(\mathbb{R},dx)}\right\|_{L^{r}(\mathbb{R},dy)}\lesssim\|\pi^{3}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{q}(\mathbb{R}\otimes\mathbb{R})}\lesssim\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})},$
and therefore when $1<r<\infty$, we have
$\|\pi^{3}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{q}(\mathbb{R}\otimes\mathbb{R})}\simeq\|g\|_{L^{r}(\mathbb{R}^{2})}.$
Proof.
First we treat (11), and we begin by proving the lower bound for the operator norm. So normalize $g$, and assume that $\|\pi^{3}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}=1$. Then fix $J$ and take an arbitrary function of the form
$f(x,y)=f_{J}(x)h_{J}(y),$
to which applying the operator yields
$\pi^{3}_{g}(f)=\sum_{I\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}\left\langle g_{I}\right\rangle_{J}h_{I}(x)h_{J}(y).$
Therefore, boundedness of $\pi^{3}_{g}$, implies that the family of one-parameter paraproducts
(13) $\pi_{g^{\prime}}(b)=\sum_{I\in\mathcal{D}}\left\langle b\right\rangle_{I}\left\langle g_{I}\right\rangle_{J}h_{I},\quad g^{\prime}=\sum_{I\in\mathcal{D}}\left\langle g_{I}\right\rangle_{J}h_{I},$
are uniformly bounded in $J$. So we have
$\sum_{I^{\prime}\subseteq I}\left\langle g_{I^{\prime}}\right\rangle_{J}^{2}\lesssim|I|,\quad I\in\mathcal{D},$
which after using Lebesgue’s differentiation theorem implies
$\sum_{I^{\prime}\subseteq I}g_{I^{\prime}}(y)^{2}\lesssim|I|,\quad I\in\mathcal{D}\quad\text{a.e}\quad y,$
as claimed in (11).
Next, we prove the upper bound for $\pi^{3}_{g}$, and to this aim again normalize $g$, such that
$\sum_{I^{\prime}\subseteq I}g_{I^{\prime}}(y)^{2}\leq|I|,\quad I\in\mathcal{D}\quad\text{a.e}\quad y,$
implying the uniform $L^{2}(\mathbb{R})$-boundedness of operators in (13), which means that for an arbitrary function of two variables $f$, and each $J$ we have
$\sum_{I\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}^{2}\left\langle g_{I}\right\rangle_{J}^{2}\lesssim\|f_{J}\|_{L^{2}(\mathbb{R})}^{2}.$
Summing over $J$ gives
$\|\pi^{3}_{g}(f)\|_{L^{2}(\mathbb{R}^{2})}^{2}=\sum_{I,J\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}^{2}\left\langle g_{I}\right\rangle_{J}^{2}\lesssim\sum_{J\in\mathcal{D}}\|f_{J}\|_{L^{2}(\mathbb{R})}^{2}=\|f\|_{L^{2}(\mathbb{R}^{2})}^{2},$
establishing $L^{2}(\mathbb{R}^{2})$-boundedness of $\pi^{3}_{g}$. Next we note that $\pi^{3}_{g}$ is a local operator simply because for any dyadic rectangle $R$, and any two dyadic intervals $I,J$, $\left\langle h_{R},\bar{\chi}_{I}\otimes h_{J}\right\rangle\neq 0$, only if $I\times J\subset R$. Therefore, according to Lemma 3.1 it is enough to show that for any $1<q<s<\infty$, and any $L^{s}$-atom $f$ with support on $\Omega$, $\pi^{3}_{g}(f)$ is an $L^{q}$-atom. For simplicity assume that $|\Omega|=1$, and for each $I$, and $y$ let
$\Omega_{I}=\{y:I\times\{y\}\subseteq\Omega\},\quad\Omega_{y}=\Omega\cap\mathbb{R}\times\{y\}.$
Next, let $\frac{1}{q}=\frac{1}{s}+\frac{1}{t}$ and note that
$S(\pi^{3}_{g}(f))(x,y)=\big(\sum_{I\times J\subseteq\Omega}\left\langle f_{J}\right\rangle_{I}^{2}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\bar{\chi}_{J}(y)\big)^{\frac{1}{2}}\leq$
$\big(\sum_{J\in\mathcal{D}}M(f_{J})^{2}(x)\bar{\chi}_{J}(y)\big)^{\frac{1}{2}}\big(\sum_{I\in\mathcal{D}}M(g_{I}\chi_{\Omega_{I}})^{2}(y)\bar{\chi}_{I}(x)\big)^{\frac{1}{2}}.$
Then Hölder’s inequality combined with $L^{s}(\mathbb{R}^{2})$-boundedness of $S_{2}M_{1}$ gives us
$\|S(\pi^{3}_{g}(f))\|_{L^{q}(\mathbb{R}^{2})}^{t}\lesssim\iint\big(\sum_{I\in\mathcal{D}}M(g_{I}(y)\chi_{\Omega_{I}})^{2}(y)\bar{\chi}_{I}(x)\big)^{\frac{t}{2}}dydx.$
Now fix $x$, and since $t>1$, we may apply Fefferman-Stein inequality and get
$\|S(\pi^{3}_{g}(f))\|_{L^{q}(\mathbb{R}^{2})}^{t}\lesssim\iint\big(\sum_{I\in\mathcal{D}}g_{I}(y)^{2}\chi_{\Omega_{I}}(y)\bar{\chi}_{I}(x)\big)^{\frac{t}{2}}dydx,$
which combined with Fubini again implies
$\|S(\pi^{3}_{g}(f))\|_{L^{q}(\mathbb{R}^{2})}^{t}\lesssim\int dy\int\big(\sum_{I\subseteq\Omega_{y}}g_{I}(y)^{2}\bar{\chi}_{I}(x)\big)^{\frac{t}{2}}dx.$
Finally, we use our assumption on $g$, and estimate the inner integral by the John-Nirenberg lemma and obtain
$\|S(\pi^{3}_{g}(f))\|_{L^{q}(\mathbb{R}^{2})}^{t}\lesssim\int|\Omega_{y}|dy=|\Omega|=1,$
which completes the proof of (11).
Now we turn to (12). Again normalize $g$ such that $\|\pi^{3}_{g}\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})\to H_{d}^{q}(\mathbb{R}\otimes\mathbb{R})}=1$, then our task is to show that
$\left\|\|g(x,y)\|_{H_{d}^{r}(\mathbb{R},dx)}\right\|_{L^{r}(\mathbb{R},dy)}^{r}\simeq\iint\big(\sum_{I\in\mathcal{D}}g_{I}(y)^{2}\bar{\chi}_{I}(x)\big)^{\frac{r}{2}}dxdy\lesssim 1.$
To this aim, for a fixed $J$ consider the operator $\pi_{g^{\prime}}$ defined in (13) and note that since
$\|\pi_{g^{\prime}}\|_{H_{d}^{p}(\mathbb{R})\to H_{d}^{q}(\mathbb{R})}\simeq\|g^{\prime}\|_{H_{d}^{r}(\mathbb{R})},$
we may find a single variable function $f_{J}$ such that
$\|f_{J}\|_{H_{d}^{p}(\mathbb{R})}^{p}=\|g^{\prime}\|_{H_{d}^{r}(\mathbb{R})}^{r},\quad\|g^{\prime}\|_{H_{d}^{r}(\mathbb{R})}\lesssim\frac{\|\pi_{g^{\prime}}(f_{J})\|_{H_{d}^{q}(\mathbb{R})}}{\|f_{J}\|_{H_{d}^{p}(\mathbb{R})}},$
implying that
$\|g^{\prime}\|_{H_{d}^{r}(\mathbb{R})}^{r}\lesssim\|\pi_{g^{\prime}}(f_{J})\|_{H_{d}^{q}(\mathbb{R})}^{q}.$
Therefore, for each $J$ we must have
(14) $\int M(f_{J})^{p}(x)dx\lesssim\int\big(\sum_{I\in\mathcal{D}}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\big)^{\frac{r}{2}}dx\lesssim\int\big(\sum_{I\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}^{2}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\big)^{\frac{q}{2}}dx.$
Now, multiply sides of the above inequality to $\chi_{J}(y)$, sum over an arbitrary finite disjoint collection of intervals $\mathcal{C}$, and integrate in $y$ to get
(15) $\iint\sum_{J\in\mathcal{C}}M(f_{J})^{p}(x)\chi_{J}(y)dxdy\lesssim\int\sum_{J\in\mathcal{C}}\big(\sum_{I\in\mathcal{D}}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\big)^{\frac{r}{2}}\chi_{J}(y)dxdy$
(16) $\lesssim\iint\sum_{J\in\mathcal{C}}\big(\sum_{I\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}^{2}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\big)^{\frac{q}{2}}\chi_{J}(y)dxdy.$
Then note that since $\mathcal{C}$ is disjoint we may write the last integrand as
$\sum_{J\in\mathcal{C}}\big(\sum_{I\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}^{2}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\big)^{\frac{q}{2}}\chi_{J}(y)=\big(\sum_{J\in\mathcal{C}}\sum_{I\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}^{2}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\chi_{J}(y)\big)^{\frac{q}{2}},$
or equivalently
$\sum_{J\in\mathcal{C}}\big(\sum_{I\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}^{2}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\big)^{\frac{q}{2}}\chi_{J}(y)=S(\pi^{3}_{g}(f))^{q}(x,y),\quad f(x,y)=\sum_{J\in\mathcal{C}}|J|^{\frac{1}{2}}f_{J}(x)h_{J}(y).$
And for the same reason, may write the first integrand in (15) as
$\sum_{J\in\mathcal{C}}M(f_{J})^{p}(x)\chi_{J}(y)=\big(\sum_{J\in\mathcal{C}}M(f_{J})^{2}(x)\chi_{J}(y)\big)^{\frac{p}{2}}=S_{2}M_{1}(f)^{p}(x,y).$
Plugging these into (15), (16) and applying boundedness of $\pi^{3}_{g}$ gives us
$\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}^{p}\lesssim\iint\sum_{J\in\mathcal{C}}\big(\sum_{I\in\mathcal{D}}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\big)^{\frac{r}{2}}\chi_{J}(y)dxdy\lesssim\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}^{q},$
which implies that
$\iint\sum_{J\in\mathcal{C}}\big(\sum_{I\in\mathcal{D}}\left\langle g_{I}\right\rangle_{J}^{2}\bar{\chi}_{I}(x)\big)^{\frac{r}{2}}\chi_{J}(y)dxdy\lesssim 1.$
Now, using Fubini first, and since $\mathcal{C}$ is arbitrary, applying Lebesgue’s differentiation theorem together with Fatou’s lemma yields
$\iint\big(\sum_{I\in\mathcal{D}}g_{I}(y)^{2}\bar{\chi}_{I}(x)\big)^{\frac{r}{2}}dydx\lesssim 1,$
which completes the proof of (12), and the theorem. ∎
Remark 3.3.
A similar argument to the one used above shows that the mixed norm in (11) is stronger than the product BMO norm.
3.2. The Operator $\pi^{4}_{g}$
Finally, we turn to the last operator, $\pi^{4}_{g}$, defined by
$\pi^{4}_{g}(f)=\sum_{I,J\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}g_{I\times J}h_{I}\otimes\bar{\chi}_{J},$
which makes it very different from the previous ones. The reason is that, unlike the other operators, $\pi^{4}_{g}$ and its adjoint are given as an expansion in an overdetermined system rather than a basis for $L^{2}(\mathbb{R}^{2})$. Therefore, one cannot directly estimate the square function of $\pi^{4}_{g}(f)$, and has to appeal to duality. Meaning that we must take $f^{\prime}$, pair it with $\pi^{4}_{g}(f)$, and note that
$\left\langle\pi^{4}_{g}(f),f^{\prime}\right\rangle=\sum_{I,J\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}\left\langle f^{\prime}_{I}\right\rangle_{J}g_{I\times J}=\left\langle\pi^{3}_{f^{\prime}}(f),g\right\rangle,$
and therefore, from boundedness of $\pi^{3}_{f^{\prime}}$ and duality it follows that
(17) $\|\pi^{4}_{g}\|_{L^{p}(\mathbb{R}^{2})\to L^{p}(\mathbb{R}^{2})}\lesssim\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})},\quad 1<p<\infty,$
(18) $\|\pi^{4}_{g}\|_{L^{p}(\mathbb{R}^{2})\to L^{q}(\mathbb{R}^{2})}\lesssim\|g\|_{L^{r}(\mathbb{R}^{2})},\quad\frac{1}{q}=\frac{1}{p}+\frac{1}{r},\quad 1<p,r,q<\infty.$
However, there is another approach, which we explain in the diagonal case. So, fix $1<p<\infty$, and for a function $g$ with finite Haar support, let
(19) $\|g\|_{p}:=\|\pi^{4}_{g}\|_{L^{p}(\mathbb{R}^{2})\to L^{p}(\mathbb{R}^{2})},$
which is identical to
$\|g\|_{p}=\sup_{\begin{subarray}{c}\|f\|_{L^{p}(\mathbb{R}^{2})}=1\\ \|f^{\prime}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}=1\end{subarray}}\Big|\sum_{I,J\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}\left\langle f^{\prime}_{I}\right\rangle_{J}g_{I\times J}\Big|.$
Then, write the sum as
(20) $\sum_{I,J\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}\left\langle f^{\prime}_{I}\right\rangle_{J}g_{I\times J}=\iint\sum_{I,J\in\mathcal{D}}g_{I\times J}|I\times J|^{-\frac{1}{2}}f_{J}(x)|J|^{-\frac{1}{2}}f^{\prime}_{I}(y)|I|^{-\frac{1}{2}}\chi_{I}(x)\chi_{J}(y)\,dxdy,$
and let $G(x,y)$ be the $\mathcal{D}$ by $\mathcal{D}$” matrix defined by
(21) $G(x,y)_{I,J}:=g_{I\times J}|I\times J|^{-\frac{1}{2}}\chi_{I}(x)\chi_{J}(y),\quad I,J\in\mathcal{D},$
and similarly define the two vectors in $l^{2}(\mathcal{D})$ by
$\vec{f}(x,y)_{J}:=f_{J}(x)|J|^{-\frac{1}{2}}\chi_{J}(x),\quad J\in\mathcal{D},\quad\vec{f^{\prime}}(x,y)_{I}:=f^{\prime}_{I}(y)|I|^{-\frac{1}{2}}\chi_{I}(x),\quad I\in\mathcal{D}.$
Then note that the integrand in (20) can be written as
$\sum_{I,J\in\mathcal{D}}G(x,y)_{I,J}\vec{f}(x,y)_{J}\vec{f^{\prime}}(x,y)_{I}=\left\langle G(x,y)\vec{f}(x,y),\vec{f^{\prime}}(x,y)\right\rangle,$
where $G(x,y)\vec{f}(x,y)$ is understood as the multiplication of a matrix with a vector, and the inner product is in $l^{2}(\mathcal{D})$. From this point of view, we may bound the right-hand side of (20) by
$\Big|\sum_{I,J\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}\left\langle f^{\prime}_{I}\right\rangle_{J}g_{I\times J}\Big|\leq\iint\big|\left\langle G(x,y)\vec{f}(x,y),\vec{f^{\prime}}(x,y)\right\rangle\big|\,dxdy,$
which implies that
$\Big|\sum_{I,J\in\mathcal{D}}\left\langle f_{J}\right\rangle_{I}\left\langle f^{\prime}_{I}\right\rangle_{J}g_{I\times J}\Big|$ $\leq\iint\|G(x,y)\|_{l^{2}(\mathcal{D})\to l^{2}(\mathcal{D})}\|\vec{f}(x,y)\|_{l^{2}(\mathcal{D})}\|\vec{f^{\prime}}(x,y)\|_{l^{2}(\mathcal{D})}\,dxdy$
$\leq\Big\|\|G(\cdot)\|_{l^{2}(\mathcal{D})\to l^{2}(\mathcal{D})}\Big\|_{L^{\infty}(\mathbb{R}^{2})}\iint\|\vec{f}(x,y)\|_{l^{2}(\mathcal{D})}\|\vec{f^{\prime}}(x,y)\|_{l^{2}(\mathcal{D})}\,dxdy$
$\leq\Big\|\|G(\cdot)\|_{l^{2}(\mathcal{D})\to l^{2}(\mathcal{D})}\Big\|_{L^{\infty}(\mathbb{R}^{2})}\Big\|\|\vec{f}(\cdot)\|_{l^{2}(\mathcal{D})}\Big\|_{L^{p}(\mathbb{R}^{2})}\Big\|\|\vec{f^{\prime}}(\cdot)\|_{l^{2}(\mathcal{D})}\Big\|_{L^{p^{\prime}}(\mathbb{R}^{2})}.$
Then note that, since
$\Big\|\|\vec{f}(\cdot)\|_{l^{2}(\mathcal{D})}\Big\|_{L^{p}(\mathbb{R}^{2})}\simeq\|f\|_{L^{p}(\mathbb{R}^{2})},\quad\Big\|\|\vec{f^{\prime}}(\cdot)\|_{l^{2}(\mathcal{D})}\Big\|_{L^{p^{\prime}}(\mathbb{R}^{2})}\simeq\|f^{\prime}\|_{L^{p^{\prime}}(\mathbb{R}^{2})},$
we obtain
(22) $\|g\|_{p}\lesssim\Big\|\|G(\cdot)\|_{l^{2}(\mathcal{D})\to l^{2}(\mathcal{D})}\Big\|_{L^{\infty}(\mathbb{R}^{2})},$
a new bound for $\|g\|_{p}$.
Now, if $g$ has a tensorial form $g=b\otimes c$, then we simply have
$\|G(x,y)\|_{l^{2}(\mathcal{D})\to l^{2}(\mathcal{D})}=\big(\sum_{I\in\mathcal{D}}b_{I}^{2}\bar{\chi}_{I}(x)\big)^{\frac{1}{2}}\big(\sum_{J\in\mathcal{D}}c_{J}^{2}\bar{\chi}_{J}(y)\big)^{\frac{1}{2}}=S(b)(x)S(c)(y)=S(g)(x,y).$
Also, because of the tensorial structure of the operator we get that
$\|g\|_{p}\simeq\|b\|_{BMO_{d}(\mathbb{R})}\|c\|_{BMO_{d}(\mathbb{R})}=\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})},$
which can be much smaller than $\|S(g)\|_{L^{\infty}(\mathbb{R}^{2})}$, and thus the bound in (22) is not sharp. Nevertheless, as the next example shows, sometimes this bound is much better than the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm.
Example. For a dyadic interval $I$ in $[0,1]$ with $|I|=2^{-m}$ let, $i(I)=m$. Now fix a large number $n$, let
$g^{\prime}_{I\times J}=\begin{cases}|I\times J|^{\frac{1}{2}}&0\leq i(I),i(J)<n,\\ 0&\text{otherwise},\end{cases}$
and note that since the first $n$ generations of dyadic rectangles cover the unit square $n^{2}$ times we have $\|g^{\prime}\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}\geq n$. Indeed since $g^{\prime}$ has a tensorial structure we have $\|g^{\prime}\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}=n$. Also, note that changing the signs of coefficients of $g^{\prime}$ does not change the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm of the function. Next, take an $n\times n$ matrix $\{H_{ij}:0\leq i,j<n\}$, with $\pm 1$ entries and with small norm. A perfect example is a Hadamard matrix, which is simply an orthogonal matrix with $\pm 1$ entries and therefore its operator norm on $\mathbb{R}^{n}$ equipped with the Euclidean norm is exactly $\sqrt{n}$. Such a matrix exists when $n$ is a power of 2. Now we modify $g^{\prime}$, obtain a new function $g$ defined by
$g_{I\times J}=H_{i(I),i(J)}g^{\prime}_{I\times J},\quad I,J\in\mathcal{D},$
and note that for each $(x,y)$, the matrix $G(x,y)$ is a copy of the matrix $H$. Therefore,
$\|G(x,y)\|_{l^{2}(\mathcal{D})\to l^{2}(\mathcal{D})}=\sqrt{n},$
which combined with (22) implies
$\|g\|_{p}\lesssim\sqrt{n},$
showing that $\|g\|_{p}$ is much smaller than $\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}$.
Interestingly, this example shows that the rectangular Haar basis is an unconditional basis for the space of functions equipped with the $\|.\|_{p}$-norm. So, one may wonder whether the remedy is to consider a stronger norm defined as
$\|g\|_{p}^{\prime}:=\sup_{\begin{subarray}{c}\|f\|_{L^{p}(\mathbb{R}^{2})}=1\\ \|f^{\prime}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}=1\end{subarray}}\sum_{I,J\in\mathcal{D}}\left\langle|f_{J}|\right\rangle_{I}\left\langle|f^{\prime}_{I}|\right\rangle_{J}|g_{I\times J}|.$
However, a modification of the above example shows that, again, this quantity can be much smaller than the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm. To see this, simply let $H$ be the identity matrix, implying that $g_{I\times J}=1$ on the first $n$ generations of dyadic squares in $0,1]^{2}$. Since they cover the unit square $n$ times, we must have $\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}\geq\sqrt{n}$. But now, at each point, the operator norm of the matrix $G(x,y)$ is $1$, and thus $\|g\|_{p}^{\prime}\lesssim 1$, again a quantity much smaller than the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm. Unfortunately, we were unable to combine this new argument with the previous ones invoking the atomic decomposition, and could not obtain an improved product BMO-type norm for $\pi^{4}_{g}$.
It is notable that sometimes the norm defined in (19) is independent of $p$. First, note that although the operator $\pi^{4}_{g}$ destroys cancellation of atoms, it does not enlarge their support. Therefore, Lemma 3.1 implies that
$\|g\|_{p}\lesssim\|g\|_{q},\quad 1<p\leq q<\infty.$
Then, under the condition that
$g_{I\times J}=g_{J\times I},\quad I,J\in\mathcal{D},$
there holds
$\|g\|_{p}\simeq\|g\|_{q},\quad 1<p,q<\infty.$
To see this, let $\tilde{f}(x,y)=f(y,x)$, and then we have
$(\pi^{4}_{g})^{t}(\tilde{f})(y,x)=\pi^{4}_{g}(f)(x,y),$
which combined with duality proves the above claim.
We conclude this note by mentioning a standard application of the lower bounds in weak factorization theorems for Hardy spaces. As an example, the bound
$\|\pi^{3}_{g}\|_{L^{p}(\mathbb{R}^{2})\to L^{q}(\mathbb{R}^{2})}\simeq\|g\|_{L^{r}(\mathbb{R}^{2})},\quad\frac{1}{q}=\frac{1}{p}+\frac{1}{r},\quad 1<p,r,q<\infty,$
can be viewed as
$\|g\|_{L^{r}(\mathbb{R}^{2})}\simeq\sup_{\begin{subarray}{c}\|f\|_{L^{p}(\mathbb{R}^{2})}=1\\ \|f^{\prime}\|_{L^{q^{\prime}}(\mathbb{R}^{2})}=1\end{subarray}}|\left\langle\pi^{4}_{f^{\prime}}(f),g\right\rangle|,$
and thus it follows directly from the general atomic decomposition in [8] that, for any function $g^{\prime}\in L^{r^{\prime}}(\mathbb{R}^{2})$, there exists a sequence of positive numbers $\{\lambda_{i}\}_{i\geq 0}$, and pairs of functions $\{f_{i},f^{\prime}_{i}\}_{i\geq 0}$ in the unit balls of $L^{p}(\mathbb{R}^{2})$ and $L^{q^{\prime}}(\mathbb{R}^{2})$, such that
$g^{\prime}=\sum_{i}\lambda_{i}\pi^{4}_{f^{\prime}_{i}}(f_{i}),\quad\|g^{\prime}\|_{L^{r^{\prime}}(\mathbb{R}^{2})}\simeq\sum_{i}\lambda_{i}.$
See [13] for similar applications of this result. On the other hand, from the above example,
$\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}\not\simeq\sup_{\begin{subarray}{c}\|f\|_{L^{p}(\mathbb{R}^{2})}=1\\ \|f^{\prime}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}=1\end{subarray}}|\left\langle\pi^{3}_{f^{\prime}}(f),g\right\rangle|,$
and thus there exists a function $g^{\prime}\in H_{d}^{1}(\mathbb{R}\otimes\mathbb{R})$ for which any factorization of the form
$g^{\prime}=\sum_{i}\lambda_{i}\pi^{3}_{f^{\prime}_{i}}(f_{i}),\quad\|g^{\prime}\|_{H_{d}^{1}(\mathbb{R}\otimes\mathbb{R})}\simeq\sum_{i}\lambda_{i},\quad\|f_{i}\|_{L^{p}(\mathbb{R}^{2})}=1,\quad\|f^{\prime}_{i}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}=1,$
would fail to hold. In particular, the Haar coefficients of $g^{\prime}$ cannot be of the form
$g^{\prime}_{I\times J}=\left\langle f_{J}\otimes f^{\prime}_{I}\right\rangle_{I\times J},\quad I,J\in\mathcal{D},\quad\sum_{I,J\in\mathcal{D}}\|f_{J}\otimes f^{\prime}_{I}\|_{L^{2}(\mathbb{R}^{2})}^{2}=1.$